OpenStudy (anonymous):

A survey of 350 customers was taken a Regal Cinemas in Austin, Texas, regarding the type of movies customers liked. The following information was detemined. 196 liked drama 153 liked comedies 88 liked science fiction 59 liked drama and comedies 37 liked drama and science fiction 32 liked comedies and science fiction 21 liked all three types of movies of the customers surveyed, how many liked A. None of these types of movies? B. Only drama? C. Exactly one of these types of movies?

OpenStudy (anonymous):

a)350-83-121-40-38-16-11-21=20 b)196-38-16-21=121 c)40+121+83=244